Shift Summary for day shift, 12/15/99

Ricardo_Barbosa (rbarbosa@fnal.gov)
Wed, 15 Dec 1999 15:53:13 -0600

Shift Summary for owl shift, 12/15/99

On Shift: J. Shields, S. Taegar

Hours of beam available at all intensities: 4.5
Hours of beam available at nominal intensity: 4.5
Hours of E799 data (special runs not included): 4.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
kfrends, swapping tapes and no beam....
========================================================================
Runs ended this shift:
Run 15220 - 7.87M - BEAM_799
Run 15221 - 1.39M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

UPA, UPB & UPC 223-228
UPS 075-076

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: Some ocasional problems in this shift but nothing
scared. After noon we have been without beam...

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15220 07:10 Red Button ENGTOT - EW. First panic at mornig.

07:49 Problems with some channels at CsI: 981 (-5.2%
in laser)
and 1096 (-7.7% in laser)

98 08:15 2 channels drop out of laser illumination plot.
Julie advice to reset pipeline.

08:16 Problems with DCOR. Inneficiency equal 100%

10:15 ktev1 is gone... rebooting...

10:15 HV out of range for Mu2 and Mu3X.
Stop run 15220

15221 10:46 Start run 15221

11:16 Lost a few spills to TeV RF problems

11:23 HV for muon system is out of range... in MU3X as
well as
MU2, where there is a known problem.

11:44 Quench! No BEAM

11:58 Roger Zimmerman calls from MCR They begin lockdown of
beam stop and when recovery from quench is done will do the
150 GeV work (KAMI test).

12:00 Stop 15221.

13:45 Crate 8 back in readout for debugging

14:05 Started artificial spill.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

- ET voltages occasionally low on ET(1) in SDAQ. Ignore; we probably
need
to fix the monitoring;
- Trigger crate 8 was out of readout all nigth. Needs fixed. Many
monitoring histos wrong because of this. Still can't reproduce error;
- we have a few low-leve CsI probs: 166, 961 have comparator errors.
The
channels 981, 1096, 1524, have gain drifts. E/p a little wide in Ke3's;
- Need to label tapes;
- 150 GeV test goes till 8PM, when Run 2 tests start. For 150 GeV, leave
beam
stops in E799 swepper settings, run COSMICs. If beam nears to target,
try running muons, slowly ramp down sweepers while watching DC currents
(all the time).
Leave beam stops closed closed closed!!!!
- ACC 90(4): there is some work planned at next in-hall access.
See page 70 of loogbook.

-- 
+---------------------------------------+
|					|
|	Ricardo Franco Barbosa		|
|		  diz			|
|		Ate' Logo!		|
|					|
+---------------------------------------+